I need help taking any decimal and turning it into a fraction. Is there a formula for this? I know the easy ones like .33 is 1/3 but I need to know how to do the more complicated conversions. Thank you so much for your help!
Found 2 solutions by ReadingBoosters, josmiceli:
Answer by ReadingBoosters(3246)

When taking a decimal to a fraction, find the place value of the last digit. Then write the decimal over that place value, as the denominator.
Examples:
.27 = (27/100)
.579 = (579/1000)
.8 = (8/10)

Answer by josmiceli(19441)

Suppose you have 1/9 which is
0.1111111111 . . . to infinity
If you are just given this repeating decimal, how
do you convert it back to 1/9?
--------------------------
First, say
x = .1111111 . . .
and
10x = 1.11111 . . .
--------------------
Now do this subtraction:
10x = 1.11111 . . .
-x = .1111111 . . .
______________________
9x = 1
x = 1/9
----------
----------
Suppose you had 7/11 = .636363636 . . .
x = .6363636 . . .
100x = 63.636363 . . .
--------------------
Do the subtraction:
100x = 63.636363 . . .
-x = .6363636 . . .
99x = 63
x = 63/99
x = (7/11)x(9/9)
x = 7/11
-----------
hope this helps